OpenStudy (anonymous):

A child's sorting game has 31 shapes. Each piece is either a triangle or a square. The shapes have a total of 111 sides. Build a system of equations that can be used to find how many of each shape is used in the game.

OpenStudy (doc.brown):

Number of triangles = t Number of squares = s 3 sides for each triangle = 3t 4 sides for each square = 4s Your equations are: \[t + s = 31\]\[3t + 4s = 111\] Solve: \[t + s = 31\]\[t = 31 - s\] Substitute t in equation 2 \[3t + 4s = 111\]\[3(31 - s) + 4s = 111\]\[93 - 3s + 4s = 111\]\[4s - 3s = 111 - 93\]\[s = 18\] \[t = 31 - s\]\[t = 31 - (18)\]\[t = 13\] check \[3t + 4s = 111\]\[3(13) + 4(18) = 111\]\[39 + 72 = 111\]\[111 = 111\]